# Reasoning Beyond Limits: Advances and Open Problems for LLMs

### v1

@@ IV-K Retrieval-Augmented Generation (RAG)
Databricks Mosaic Research [75] presents a comprehensive evaluation of long-context Retrieval Augmented Generation (RAG) capabilities across several state-of-the-art language models. The authors compared models including OpenAI o1, Anthropic Claude, Google DeepMind Gemini, and Meta Llama on three diverse datasets—Databricks DocsQA, FinanceBench, and Natural Questions—by varying context lengths from 2,000 up to 2,000,000 tokens retrieved from a vector database. Using GPT-4o as the judge to assess answer correctness, the study found that Google DeepMind’s Gemini 1.5 maintains strong RAG performance even with highly extended contexts (up to 2 million tokens). In contrast, open models like Llama-3.1-405B show significant performance degradation beyond 32k tokens. OpenAI’s o1-preview model achieved the highest average performance score of 0.763 across the tested context lengths, although standard OpenAI o1 models underperformed at shorter contexts on specific datasets. Most models demonstrated peak performance in the 32k-64k token range before declining, highlighting critical insights into the optimal context lengths and trade-offs in deploying long-context RAG systems.
@@
Wang [76] offers a comprehensive investigation into Retrieval-Augmented Generation (RAG) techniques, which are widely recognized for their ability to integrate current information, reduce hallucinations, and enhance output quality in domain-specific applications. By dissecting the multi-step RAG workflow—ranging from query classification, advanced chunking, and hybrid retrieval using methods like HyDE, to effective reranking with models such as MonoT5—the study systematically evaluates various components to establish best practices that strike a balance between performance and efficiency. Additionally, the work explores multimodal retrieval strategies that improve the handling of visual inputs and accelerate the generation of multimodal content via a ”retrieval as generation” approach. Although the paper is more of a synthesis of existing methodologies rather than a breakthrough in novel techniques, its extensive experimentation and clear recommendations, including the use of Milvus for vector storage, reverse repacking strategies, and mixed-context fine-tuning for robust generation, provide valuable insights for optimizing RAG systems.
IV-K1 Chain-of-Retrieval Augmented Generation (CoRAG)
@@
Wang et al. [77] propose CoRAG (Chain-of-Retrieval Augmented Generation), a novel framework designed to enhance traditional RAG models by incorporating iterative retrieval and reasoning steps before answer generation. Conventional RAG approaches typically rely on a single retrieval phase, which can limit performance when handling complex, multi-hop queries due to incomplete or imperfect retrievals. In contrast, CoRAG dynamically reformulates the query as the model’s internal state evolves, enabling the construction of intermediate retrieval chains that better capture the necessary context. The training process leverages rejection sampling to automatically generate these chains, effectively augmenting datasets that otherwise only provide the final answer. The method employs flexible decoding strategies at inference time to manage computational demands by controlling the length and number of retrieval chains sampled. Empirical results across various benchmarks, including a notable improvement in exact match scores for multi-hop question answering and state-of-the-art performance on the KILT benchmark, validate the efficacy of CoRAG
IV-K2 RAG mechanism with ”Reason-in-Documents” module
@@
Li et al. [5] propose Search-o1, an innovative framework designed to enhance large reasoning models (LRMs) such as OpenAI-o1, which, despite their robust stepwise reasoning abilities, often encounter limitations due to insufficient knowledge during extended reasoning processes. To overcome these challenges, Search-o1 incorporates an agentic retrieval-augmented generation (RAG) mechanism that dynamically searches for and retrieves external knowledge when uncertainties arise. Furthermore, the framework includes a dedicated Reason-in-Documents module that refines and analyzes the verbose retrieved content, ensuring that only coherent and relevant information is integrated into the reasoning chain. Experimental evaluations across diverse complex reasoning tasks in science, mathematics, coding, and open-domain question answering benchmarks demonstrate that Search-o1 significantly improves performance, thereby enhancing the reliability and versatility of LRMs in handling intricate reasoning challenges.
IV-K3 Contextualized Document Embeddings
@@
Morris et al. [78] advanced the field of neural retrieval by proposing contextualized document embeddings that incorporate information from neighboring documents, addressing a key limitation of traditional dense embeddings that are generated in isolation. The authors introduce two complementary approaches: one modifies the contrastive learning objective to explicitly integrate neighbor information into the intra-batch loss, and the other develops a new architectural design that encodes context from neighboring documents directly into the representation. This dual strategy leads to significant improvements over standard biencoder methods, particularly in out-of-domain retrieval scenarios. Notably, the method achieves state-of-the-art performance on the MTEB benchmark without relying on standard techniques such as hard negative mining, score distillation, or extremely large batch sizes. By clustering similar documents to identify neighbors and seamlessly incorporating their context during both training and encoding, the approach yields more robust and context-aware embeddings. This work not only enhances retrieval performance for applications such as Retrieval-Augmented Generation (RAG) but also offers a versatile framework that can be applied to any contrastive learning dataset or biencoder model, underscoring its broad applicability and potential impact on improving retrieval systems.
IV-K4 RAG with Self-Reasoning
@@
Xia et al. [79] presents an innovative self-reasoning framework designed to enhance the reliability and traceability of Retrieval-Augmented Language Models (RALMs), addressing persistent challenges such as irrelevant document retrieval and insufficient citation in generated responses. The proposed approach leverages the internal reasoning trajectories of the LLM through a structured three-stage process. First, the Relevance-Aware Process (RAP) instructs the model to assess the relevance of retrieved documents to the query. Next, the Evidence-Aware Selective Process (EAP) extracts key sentences from these documents, along with justifications for their selection. Finally, the Trajectory Analysis Process (TAP) synthesizes these reasoning paths to produce a final answer. Evaluated on four diverse public datasets—including short-form and long-form QA as well as fact verification tasks—the framework demonstrates a performance improvement of over 10% compared to standard RAG approaches when applied to a Llama 2 7B model, all while requiring only 2,000 training samples generated using GPT-4. This method not only improves the factual accuracy and trustworthiness of RALMs but also offers a scalable solution for integrating structured, self-generated reasoning into retrieval-based systems.
IV-L Evolutionary search strategy
@@
Lee et al. [87] introduced Mind Evolution, an innovative evolutionary search strategy designed to enhance inference time computation in LLMs for natural language planning tasks. Rather than requiring an explicit formalization of the inference problem, this approach leverages a language model to iteratively generate, evaluate, and refine candidate solutions, using a programmatic evaluator to guide the process. Mind Evolution significantly outperforms traditional inference strategies such as Best-of-N and Sequential Revision by systematically recombining and optimizing candidate responses over multiple generations. Experimental results on benchmarks like TravelPlanner and Natural Plan demonstrate a success rate exceeding 98% when using Gemini 1.5 Pro, all while maintaining a more cost-effective inference process compared to sequential revision methods. Additionally, the framework supports a two-stage approach—initially employing a faster model for simpler tasks, then transitioning to a more robust model for complex cases—thereby showcasing its adaptability and efficiency in tackling planning problems that are challenging to formalize but straightforward to evaluate.
IV-M Data leakage and deduplication
@@
SemHash [88] is a high-performance semantic deduplication library designed to address critical challenges in data leakage and deduplication during the training of LLMs. SemHash leverages Model2Vec embeddings in conjunction with an approximate nearest neighbor (ANN) similarity search via Vicinity, enabling rapid and efficient deduplication of large-scale text datasets. Notably, the system can process 1.8 million WikiText records in just 83 seconds on a CPU, highlighting its remarkable speed and scalability. By focusing on semantic similarity rather than strict exact matching, SemHash can identify near-duplicate records, thereby enhancing the quality of both training and evaluation datasets—whether for cleaning single datasets or for preventing cross-dataset leakage between training and testing splits. The library offers a straightforward Python API with minimal dependencies. It supports the integration of custom encoders, such as those provided by sentence-transformers, along with built-in tools for inspecting duplicates and adjusting similarity thresholds. Benchmark results across 17 popular datasets underscore its effectiveness, making SemHash a valuable tool for researchers and practitioners working on preparing high-quality datasets for LLM training.
IV-N A non-agent framework
@@
Xia et al. [89] challenges the prevailing reliance on complex autonomous software agents for automating software development tasks by introducing Agentless, a streamlined, non-agent framework that simplifies the repair process into distinct phases of localization and repair. Instead of employing elaborate agent libraries and decision-making for tool usage, Agentless leverages a tree-structured repository view alongside embedding-based retrieval techniques to isolate suspicious files and extract function and class signatures—the “skeletons”—of code under investigation. The framework then generates targeted search/replace diffs and corresponding reproduction tests, ultimately validating the best patch via majority voting and regression testing. Evaluated on the SWE-bench Lite benchmark, Agentless not only achieves the highest performance at a 27.33% solve rate but also incurs significantly lower costs (\$0.34 per issue) compared to existing agent-based approaches. These findings underscore that a focused, interpretable pipeline can effectively address real-world software engineering problems while avoiding the complexity and pitfalls of more verbose autonomous agent systems, thereby resetting the baseline and inspiring future work in autonomous software development.
IV-O Chain-of-Thought (CoT)
@@
Chain-of-thought prompting [9] guides a language model to articulate intermediate reasoning steps before reaching a final answer. This method mirrors human problem-solving by decomposing complex tasks into simpler, sequential parts. It not only boosts the model’s accuracy on challenging tasks—such as multi-step arithmetic and symbolic reasoning—but also makes its internal decision-making more transparent, thereby aiding in error analysis and debugging.
@@
Cui et al. [121] investigated the effectiveness of extending few-shot chain-of-thought (CoT) prompting by incorporating both correct and incorrect reasoning paths along with detailed explanations for the errors. Traditional CoT prompting methods typically provide a sequence of in-context learning steps (Stepwise ICL) that guide the transformer’s reasoning process; however, the authors theoretically demonstrate that integrating coherent reasoning from earlier steps enhances error correction and improves prediction accuracy. Their analysis reveals that transformers are particularly sensitive to mistakes made in intermediate reasoning steps rather than solely focusing on the final outcome. Building on this insight, the proposed approach augments standard few-shot CoT prompts with deliberately generated incorrect reasoning examples—accompanied by explanations that clarify the nature of the mistakes. The implementation involves selecting a base large language model (e.g., GPT-3.5, GPT-4, Gemini Pro, or DeepSeek 67B), generating correct step-by-step reasoning samples, creating and annotating incorrect reasoning paths, and combining these elements in the demonstration prompts. Experimental results indicate that using model-generated incorrect examples with corresponding explanations significantly enhances reasoning performance across diverse LLMs, while omitting the explanatory component or relying on handcrafted errors can be detrimental.
IV-O1 Chain of Continuous Thought
@@
Meta team introduces Coconut (Chain of Continuous Thought) [80], a novel reasoning paradigm for LLMs that shifts from traditional chain-of-thought (CoT) reasoning in natural language to a latent-space-based approach. Instead of decoding hidden states into textual tokens, Coconut directly feeds them back into the model as input, enabling a more flexible and efficient reasoning process. By utilizing two modes—Language Mode for standard text generation and Latent Mode for iterative hidden-state processing—Coconut facilitates a chain of continuous thoughts, allowing the model to explore multiple alternative reasoning paths simultaneously. This method, marked by special ¡bot¿ and ¡eot¿ tokens, enables breadth-first search (BFS)-like reasoning, which outperforms CoT in complex planning tasks such as ProntoQA and ProsQA, particularly those requiring substantial backtracking. Additionally, Coconut significantly reduces token generation during inference, making it more efficient while maintaining accuracy. Inspired by iterative CoT (iCoT), the model is trained using a multi-stage curriculum learning strategy that gradually replaces language-based reasoning with latent-space reasoning, ensuring a smooth transition. The findings demonstrate that leveraging latent representations enhances reasoning efficiency, making Coconut a promising advancement in LLM-based problem-solving and strategic planning.
IV-O2 Chain-of-Retrieval Augmented Generation (CoRAG)
@@
Wang et al. [77] introduced Chain-of-Retrieval Augmented Generation (CoRAG), a novel approach designed to address a key limitation in traditional RAG systems—namely, the reliance on a single retrieval step that often falls short when handling complex, multi-hop queries. Conventional RAG methods retrieve supporting information once before the generation phase, which can lead to incomplete or suboptimal evidence for intricate queries. In contrast, CoRAG enables dynamic query reformulation by allowing the model to iteratively retrieve and integrate information, effectively building a chain of reasoning that better supports the final answer. To overcome the challenge of limited annotated data for intermediate steps, the authors employ rejection sampling to automatically generate plausible retrieval chains from existing RAG datasets that typically only include the final correct answer. During inference, they propose flexible decoding strategies that adjust the length and number of retrieval chains, providing a trade-off between computational cost and accuracy. Experimental results demonstrate that CoRAG significantly outperforms strong baselines—improving exact match scores by over 10 points on multi-hop question answering tasks—and sets a new state-of-the-art on the KILT benchmark across various knowledge-intensive tasks. Comprehensive analyses of the model’s scaling behavior further underscore its potential to guide future research toward developing more factual and grounded foundation models.
@@
IV-O3 Monte Carlo Tree Search (OmegaPRM)

### v2

@@ VI-K Retrieval-Augmented Generation (RAG)
Databricks Mosaic Research [databricks2023longcontext] presents a comprehensive evaluation of long-context Retrieval Augmented Generation (RAG) capabilities across several state-of-the-art language models. The authors compared models including OpenAI o1, Anthropic Claude, Google DeepMind Gemini, and Meta Llama on three diverse datasets Databricks DocsQA, FinanceBench, and Natural Questions by varying context lengths from 2,000 up to 2,000,000 tokens retrieved from a vector database. Using GPT-4o as the judge to assess answer correctness, the study found that Google DeepMind’s Gemini 1.5 maintains strong RAG performance even with highly extended contexts (up to 2 million tokens). In contrast, open models like Llama-3.1-405B show significant performance degradation beyond 32k tokens. OpenAI’s o1-preview model achieved the highest average performance score of 0.763 across the tested context lengths, although standard OpenAI o1 models underperformed at shorter contexts on specific datasets. Most models demonstrated peak performance in the 32k-64k token range before declining, highlighting critical insights into the optimal context lengths and trade-offs in deploying long-context RAG systems.
@@
Wang [wang2024searching] offers a comprehensive investigation into Retrieval-Augmented Generation (RAG) techniques, which are widely recognized for their ability to integrate current information, reduce hallucinations, and enhance output quality in domain-specific applications. By dissecting the multi-step RAG workflow ranging from query classification, advanced chunking, and hybrid retrieval using methods like HyDE, to effective reranking with models such as MonoT5 the study systematically evaluates various components to establish best practices that strike a balance between performance and efficiency. Additionally, the work explores multimodal retrieval strategies that improve the handling of visual inputs and accelerate the generation of multimodal content via a ”retrieval as generation” approach. Although the paper is more of a synthesis of existing methodologies rather than a breakthrough in novel techniques, its extensive experimentation and clear recommendations, including the use of Milvus for vector storage, reverse repacking strategies, and mixed-context fine-tuning for robust generation, provide valuable insights for optimizing RAG systems.
VI-K1 Chain-of-Retrieval Augmented Generation (CoRAG)
@@
Wang et al. [wang2025chain] propose CoRAG (Chain-of-Retrieval Augmented Generation), a novel framework designed to enhance traditional RAG models by incorporating iterative retrieval and reasoning steps before answer generation. Conventional RAG approaches typically rely on a single retrieval phase, which can limit performance when handling complex, multi-hop queries due to incomplete or imperfect retrievals. In contrast, CoRAG dynamically reformulates the query as the model’s internal state evolves, enabling the construction of intermediate retrieval chains that better capture the necessary context. The training process leverages rejection sampling to automatically generate these chains, effectively augmenting datasets that otherwise only provide the final answer. The method employs flexible decoding strategies at inference time to manage computational demands by controlling the length and number of retrieval chains sampled. Empirical results across various benchmarks, including a notable improvement in exact match scores for multi-hop question answering and state-of-the-art performance on the KILT benchmark, validate the efficacy of CoRAG
VI-K2 RAG mechanism with ”Reason-in-Documents” module
@@
Li et al. [li2025search] propose Search-o1, an innovative framework designed to enhance large reasoning models (LRMs) such as OpenAI-o1, which, despite their robust stepwise reasoning abilities, often encounter limitations due to insufficient knowledge during extended reasoning processes. To overcome these challenges, Search-o1 incorporates an agentic retrieval-augmented generation (RAG) mechanism that dynamically searches for and retrieves external knowledge when uncertainties arise. Furthermore, the framework includes a dedicated Reason-in-Documents module that refines and analyzes the verbose retrieved content, ensuring that only coherent and relevant information is integrated into the reasoning chain. Experimental evaluations across diverse complex reasoning tasks in science, mathematics, coding, and open-domain question answering benchmarks demonstrate that Search-o1 significantly improves performance, thereby enhancing the reliability and versatility of LRMs in handling intricate reasoning challenges.
VI-K3 Contextualized Document Embeddings
@@
Morris et al. [morris2024contextual] advanced the field of neural retrieval by proposing contextualized document embeddings that incorporate information from neighboring documents, addressing a key limitation of traditional dense embeddings that are generated in isolation. The authors introduce two complementary approaches: one modifies the contrastive learning objective to explicitly integrate neighbor information into the intra-batch loss, and the other develops a new architectural design that encodes context from neighboring documents directly into the representation. This dual strategy leads to significant improvements over standard biencoder methods, particularly in out-of-domain retrieval scenarios. Notably, the method achieves state-of-the-art performance on the MTEB benchmark without relying on standard techniques such as hard negative mining, score distillation, or extremely large batch sizes. By clustering similar documents to identify neighbors and seamlessly incorporating their context during both training and encoding, the approach yields more robust and context-aware embeddings. This work not only enhances retrieval performance for applications such as Retrieval-Augmented Generation (RAG) but also offers a versatile framework that can be applied to any contrastive learning dataset or biencoder model, underscoring its broad applicability and potential impact on improving retrieval systems.
VI-K4 RAG with Self-Reasoning
@@
Xia et al. [xia2024improving] presents an innovative self-reasoning framework designed to enhance the reliability and traceability of Retrieval-Augmented Language Models (RALMs), addressing persistent challenges such as irrelevant document retrieval and insufficient citation in generated responses. The proposed approach leverages the internal reasoning trajectories of the LLM through a structured three-stage process. First, the Relevance-Aware Process (RAP) instructs the model to assess the relevance of retrieved documents to the query. Next, the Evidence-Aware Selective Process (EAP) extracts key sentences from these documents, along with justifications for their selection. Finally, the Trajectory Analysis Process (TAP) synthesizes these reasoning paths to produce a final answer. Evaluated on four diverse public datasets including short-form and long-form QA as well as fact verification tasks the framework demonstrates a performance improvement of over 10% compared to standard RAG approaches when applied to a Llama 2 7B model, all while requiring only 2,000 training samples generated using GPT-4. This method not only improves the factual accuracy and trustworthiness of RALMs but also offers a scalable solution for integrating structured, self-generated reasoning into retrieval-based systems.
VI-L Evolutionary search strategy
@@
Lee et al. [lee2025evolving] introduced Mind Evolution, an innovative evolutionary search strategy designed to enhance inference time computation in LLMs for natural language planning tasks. Rather than requiring an explicit formalization of the inference problem, this approach leverages a language model to iteratively generate, evaluate, and refine candidate solutions, using a programmatic evaluator to guide the process. Mind Evolution significantly outperforms traditional inference strategies such as Best-of-N and Sequential Revision by systematically recombining and optimizing candidate responses over multiple generations. Experimental results on benchmarks like TravelPlanner and Natural Plan demonstrate a success rate exceeding 98% when using Gemini 1.5 Pro, all while maintaining a more cost-effective inference process compared to sequential revision methods. Additionally, the framework supports a two-stage approach initially employing a faster model for simpler tasks, then transitioning to a more robust model for complex cases thereby showcasing its adaptability and efficiency in tackling planning problems that are challenging to formalize but straightforward to evaluate.
VI-M Data leakage and deduplication
@@
SemHash [minishlab2025semhash] is a high-performance semantic deduplication library designed to address critical challenges in data leakage and deduplication during the training of LLMs. SemHash leverages Model2Vec embeddings in conjunction with an approximate nearest neighbor (ANN) similarity search via Vicinity, enabling rapid and efficient deduplication of large-scale text datasets. Notably, the system can process 1.8 million WikiText records in just 83 seconds on a CPU, highlighting its remarkable speed and scalability. By focusing on semantic similarity rather than strict exact matching, SemHash can identify near-duplicate records, thereby enhancing the quality of both training and evaluation datasets whether for cleaning single datasets or for preventing cross-dataset leakage between training and testing splits. The library offers a straightforward Python API with minimal dependencies. It supports the integration of custom encoders, such as those provided by sentence-transformers, along with built-in tools for inspecting duplicates and adjusting similarity thresholds. Benchmark results across 17 popular datasets underscore its effectiveness, making SemHash a valuable tool for researchers and practitioners working on preparing high-quality datasets for LLM training.
VI-N A non-agent framework
@@
Xia et al. [xia2024agentless] challenges the prevailing reliance on complex autonomous software agents for automating software development tasks by introducing Agentless, a streamlined, non-agent framework that simplifies the repair process into distinct phases of localization and repair. Instead of employing elaborate agent libraries and decision-making for tool usage, Agentless leverages a tree-structured repository view alongside embedding-based retrieval techniques to isolate suspicious files and extract function and class signatures the “skeletons” of code under investigation. The framework then generates targeted search/replace diffs and corresponding reproduction tests, ultimately validating the best patch via majority voting and regression testing. Evaluated on the SWE-bench Lite benchmark, Agentless not only achieves the highest performance at a 27.33% solve rate but also incurs significantly lower costs (\$0.34 per issue) compared to existing agent-based approaches. These findings underscore that a focused, interpretable pipeline can effectively address real-world software engineering problems while avoiding the complexity and pitfalls of more verbose autonomous agent systems, thereby resetting the baseline and inspiring future work in autonomous software development.
VI-O Chain-of-Thought (CoT)
@@
Chain-of-thought prompting [wei2022chain] guides a language model to articulate intermediate reasoning steps before reaching a final answer. This method mirrors human problem-solving by decomposing complex tasks into simpler, sequential parts. It not only boosts the model’s accuracy on challenging tasks such as multi-step arithmetic and symbolic reasoning but also makes its internal decision-making more transparent, thereby aiding in error analysis and debugging.
@@
Cui et al. [cui2024theoretical] investigated the effectiveness of extending few-shot chain-of-thought (CoT) prompting by incorporating both correct and incorrect reasoning paths along with detailed explanations for the errors. Traditional CoT prompting methods typically provide a sequence of in-context learning steps (Stepwise ICL) that guide the transformer’s reasoning process; however, the authors theoretically demonstrate that integrating coherent reasoning from earlier steps enhances error correction and improves prediction accuracy. Their analysis reveals that transformers are particularly sensitive to mistakes made in intermediate reasoning steps rather than solely focusing on the final outcome. Building on this insight, the proposed approach augments standard few-shot CoT prompts with deliberately generated incorrect reasoning examples accompanied by explanations that clarify the nature of the mistakes. The implementation involves selecting a base large language model (e.g., GPT-3.5, GPT-4, Gemini Pro, or DeepSeek 67B), generating correct step-by-step reasoning samples, creating and annotating incorrect reasoning paths, and combining these elements in the demonstration prompts. Experimental results indicate that using model-generated incorrect examples with corresponding explanations significantly enhances reasoning performance across diverse LLMs, while omitting the explanatory component or relying on handcrafted errors can be detrimental.
VI-O1 Chain of Continuous Thought
@@
Meta team introduces Coconut (Chain of Continuous Thought) [hao2024training], a novel reasoning paradigm for LLMs that shifts from traditional chain-of-thought (CoT) reasoning in natural language to a latent-space-based approach. Instead of decoding hidden states into textual tokens, Coconut directly feeds them back into the model as input, enabling a more flexible and efficient reasoning process. By utilizing two modes Language Mode for standard text generation and Latent Mode for iterative hidden-state processing Coconut facilitates a chain of continuous thoughts, allowing the model to explore multiple alternative reasoning paths simultaneously. This method, marked by special <bot><bot> and <eot><eot> tokens, enables breadth-first search (BFS)-like reasoning, which outperforms CoT in complex planning tasks such as ProntoQA and ProsQA, particularly those requiring substantial backtracking. Additionally, Coconut significantly reduces token generation during inference, making it more efficient while maintaining accuracy. Inspired by iterative CoT (iCoT), the model is trained using a multi-stage curriculum learning strategy that gradually replaces language-based reasoning with latent-space reasoning, ensuring a smooth transition. The findings demonstrate that leveraging latent representations enhances reasoning efficiency, making Coconut a promising advancement in LLM-based problem-solving and strategic planning.
VI-O2 Chain-of-Retrieval Augmented Generation (CoRAG)
@@
Wang et al. [wang2025chain] introduced Chain-of-Retrieval Augmented Generation (CoRAG), a novel approach designed to address a key limitation in traditional RAG systems namely, the reliance on a single retrieval step that often falls short when handling complex, multi-hop queries. Conventional RAG methods retrieve supporting information once before the generation phase, which can lead to incomplete or suboptimal evidence for intricate queries. In contrast, CoRAG enables dynamic query reformulation by allowing the model to iteratively retrieve and integrate information, effectively building a chain of reasoning that better supports the final answer. To overcome the challenge of limited annotated data for intermediate steps, the authors employ rejection sampling to automatically generate plausible retrieval chains from existing RAG datasets that typically only include the final correct answer. During inference, they propose flexible decoding strategies that adjust the length and number of retrieval chains, providing a trade-off between computational cost and accuracy. Experimental results demonstrate that CoRAG significantly outperforms strong baselines improving exact match scores by over 10 points on multi-hop question answering tasks and sets a new state-of-the-art on the KILT benchmark across various knowledge-intensive tasks. Comprehensive analyses of the model’s scaling behavior further underscore its potential to guide future research toward developing more factual and grounded foundation models.
@@
VI-O3 Monte Carlo Tree Search (OmegaPRM)